# On the Complexity of Minimizing Energy Consumption of Partitioning DAG Tasks

### v1

@@ 4 Energy-Saving Partition of DAG
Let $v$ be as stipulated in Claim 6. We consider first the case where $\Gamma^{+}_{G}{[v]}\subseteq X$. Clearly, $X^{\prime}=X\cup\{v^{+},v^{-}\}$ and $s\in X^{\prime}$ in this case. We need to prove that $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$. To this end, observe that none of $\{(u,v^{-}):u\in\Gamma^{+}_{G}{(v)}\}$ is in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$, due to the infinite weight of arcs in this set. It follows that $v^{-}\in X$. If $v^{+}\in X$, it holds that $(X^{\prime},Y^{\prime})=(X,Y)$, and thus $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))=w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$; we are done. Assume now that $v^{+}\in Y$. Then, as $\Gamma^{-}_{G^{\prime}}{(v^{+})}=\{v\}$, we have that $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))-w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))=w((v,v^{+}))\geq 0$.
We consider now the case where $\Gamma^{+}_{G}{[v]}\subseteq Y$. In this case, $X^{\prime}=Y\cup\{v^{+},v^{-}\}$, $Y^{\prime}=X\setminus\{v^{+},v^{-}\}$, and $s\in Y^{\prime}$. Observe that none of $\{(v^{+},u):u\in\Gamma^{+}_{G}{(v)}\}$ is in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$, due to the infinite weight of arcs in this set. It follows that $v^{+}\in Y$. If $v^{-}\in Y$, it holds that $(Y^{\prime},X^{\prime})=(X,Y)$, and thus $w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))=w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$; we are done. Assume now that $v^{-}\in X$. Then, as $\Gamma^{+}_{G^{\prime}}{(v^{-})}=\{v\}$, we have that $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))-w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))=w((v^{-},v))\geq 0$. ∎
By Claim 6, we may assume that, for every $v\in V$ such that $\Gamma^{+}_{G}{(v)}\neq\emptyset$, if $\Gamma^{+}_{G}{[v]}$ are contained in the same $Z\in\{X,Y\}$, then $\{v^{+},v^{-}\}$ are also contained in $Z$.
The following claims are mainly due to the infinite weight of arcs leaving $v^{+}$ and entering $v^{-}$.
@@
Let $v\in X\cap V(G)$ such that $\Gamma^{+}_{G}{(v)}\cap(Y\setminus\{t\})\neq\emptyset$. Then, $(v,v^{+})\in{{\textsf{cut}}}_{G^{\prime}}(X,Y)$ and none of the arcs entering or leaving $v^{-}$ is contained in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$.
@@
Let $v$ be as stipulated in the claim. Let $u$ be any arbitrary vertex from $\Gamma^{+}_{G}{(v)}\cap(Y\setminus\{t\})$. It must hold that $v^{+}\in Y$, since otherwise the arc $(v^{+},u)$ which has infinite weight is contained in ${{\textsf{cut}}}_{Y}(G^{\prime},X)$, contradicting that $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$. As $v$ is the only outneighbor of $v^{-}$ in $G$ and $v\in X$, we know that the only arc $(v^{-},v)$ leaving $v^{-}$ is excluded from ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$. Finally, recall that $\Gamma^{-}_{G^{\prime}}{(v^{-})}=\{u\in V(G):u\in\Gamma^{+}_{G}{(v)}\}$, and by the definition of the function $w$, every arc $(u,v^{-})$ where $u\in\Gamma^{+}_{G}{(v)}$ has infinite weight. As $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$, we know that none of the arcs entering $v^{-}$ is contained in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$. ∎
By Claim 7, for every $v\in X\cap V(G)$ having at least one outneighbor from $Y\setminus\{t\}$ in $G$, we may assume that $v^{+}\in Y$. Moreover, as $v$ is the only outneighbor of $v^{-}$ in $G^{\prime}$, we may assume that $v^{-}\in X$.
By similar reasoning, we obtain the following claim.
@@
Let $v\in Y\cap V(G)$ such that $\Gamma^{+}_{G}{(v)}\cap(X\setminus\{s\})\neq\emptyset$. Then, $(v^{-},v)\in{{\textsf{cut}}}_{G^{\prime}}(X,Y)$, and none of the arcs entering or leaving $v^{+}$ is contained in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$.
By Claim 8, for every $v\in Y\cap V(G)$ having at least one outneighbor from $X\setminus\{s\}$ in $G$, we may assume that $v^{-}\in X$ and, moreover, since $v$ is the only inneighbor of $v^{+}$ in $G^{\prime}$, we may assume that $v^{+}\in Y$.
Let $V_{1}=V(G)\cap X$ and let $V_{2}=V(G)\cap Y$. By the definition of the function $w$, we have that
@@
Let $V_{1}^{\prime}$ be the set consisting of all vertices from $V_{1}$ having at least one outneighbor from $Y$ in $G^{\prime}$. Similarly, let $V_{2}^{\prime}$ be the set of vertices of $V_{2}$ having at least one outneighbor from $X$ in $G^{\prime}$. By Claims 68, the following statements hold:
  • 1.
    For every $i\in[2]$ and every $v\in V_{i}\setminus V_{i}^{\prime}$, $v^{+}$ is contained in the same component with $v$ in the partition $(X,Y)$.
  • 2.
@@
The sum of the left sides of Equalities (9)–(11) is $\left(\sum_{v\in V_{i},i\in[2]}p(v,i)\right)+{\textsf{BTran}_{G}((V_{1},V_{2}),q)}$, and the sum of the right sides of them is $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$. From $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$, we know that the instance of ESP-DAG is a Yes-instance. ∎
As DM-$s$-$t$-Cut can be solved in time $O(n\cdot m\cdot\log(n^{2}/m))$ [34], by Theorem 2, we reach the following result.
@@
ESP-DAG with $k=2$ can be solved in time $O(n\cdot m\cdot\log(n^{2}/m))$, where $n$ and $m$ are respectively the number of vertices and the number of edges of the input graph.
Next, we derive a pseudo polynomial-time algorithm for a special case of ESP-DAG where the input DAG is a directed path. Tasks having such precedence dependencies are relevant to many applications (see, e.g., [35, 36]).
@@
ESP-DAG can be solved in time $O(k^{3}\cdot n\cdot M)$ when the given DAG is a directed path of $n$ vertices, where $M$ is the number of bits to encode any number in the range of the energy consumption function $q$ and in the range of the function $p$.
@@
Let $I=(G,p,q,r)$ be an instance of ESP-DAG, where $G$ is a directed path, $p:V(G)\times[k]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$ are two functions, and $r$ is a number. We derive a dynamic programming algorithm to solve the problem as follows. Let $(v_{1},v_{2},\dots,v_{n})$ be the directed path. We maintain a table $S(i,j)$ where $i\in[n]$ and $j\in[k]$. Specifically, $S(i,j)$ is defined as the value of an optimal solution to the instance $I$ restricted to the first $i$ vertices under the restriction that $v_{i}$ is assigned to the machine $j$. More precisely, we define
@@
By the definition, for each $j\in[k]$, we have that $S(1,j)=p(v_{1},j)$. We use the following recursion to update the table:
@@
After the table is computed, we conclude that $I$ is a Yes-instance if and only if
@@
It remains to analyze the running time of the algorithm. As there are at most $n\cdot k$ entries in the table, and computing each entry needs to check at most $k$ other entries and perform addition/comparison operations $O(k)$ times, the table can be computed in time $O(nk^{3})$. ∎
@@ 5 Sized-Bounded Energy-Saving Partition of DAG with Two Machines
Now let us switch our focus to the natural variant SB-ESBP-DAG where there are only two machines and one of them is capable of executing a limited number of tasks.
Recall that SBM-$s$-$t$-Cut is exactly Multiway Cut where $k=2$ with an additional restriction that one of the components in the desired bipartition contains at most $\ell$ vertices. Then, based on the reduction in the proof of Theorem 1 and the fact that SBM-$s$-$t$-Cut is NP-hard, we know that SB-ESBP-DAG is NP-hard as well. Unlike ESP-DAG, this variate possesses a natural parameter $\ell$, the number of tasks one of the machines is capable of handling. This motivates us to consider the parameterized complexity of SB-ESBP-DAG with respect to $\ell$.
A concise introduction to parameterized complexity theory can be beneficial for readers unfamiliar with the concept. A parameterized problem is a subset of $\Sigma^{*}\times\mathbb{N}$, where $\Sigma$ is a finite alphabet. A parameterized problem can be either fixed-parameter tractable (FPT) or fixed-parameter intractable. In particular, a parameterized problem is FPT if there is an algorithm which correctly determines for each instance $(I,\kappa)$ of the problem whether $(I,\kappa)$ is a Yes-instance in time $O(f(\kappa)\cdot|I|^{O(1)})$, where $f$ is a computable function and $|I|$ is the size of $I$. W[1]-hard problems are considered intractable, in the sense that unless FPT$=$W[1] (which is widely believed to be unlikely), they do not admit any FPT-algorithms. For greater details on parameterized complexity theory, we refer to [8, 9, 10].
We show that SB-ESBP-DAG is W[1]-hard with respect to this parameter. To this end, through a trivial reduction from Clique on regular graphs, we show that SBM-$s$-$t$-Cut is W[1]-hard with respect to $\ell$ even in a special case. Despite the widespread attention this problem has received, it is remarkable that such a W[1]-hardness result has not been previously documented. Therefore, our reduction bolsters the NP-hardness of SBM-$s$-$t$-Cut as examined in [11].
@@
SBM-$s$-$t$-Cut is W[1]-hard with respect to $\ell$. Moreover, this holds even if the edges in the given graph have at most two different weights.
@@
Let $(G,\ell)$ be an instance of Clique where $G$ is a regular graph of $n$ vertices. Let $d$ denote the degree of vertices in $G$. Let $G^{\prime}$ be the graph obtained from $G$ by adding two vertices $s$ and $t$ both of which are adjacent to all vertices of $G$. Let $w:E(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ be a function such that for every edge $\{x,y\}\in E(G^{\prime})$ it holds that
$w(\{x,y\})=\begin{cases}d\cdot\ell+\frac{\ell\cdot(\ell-1)}{2}+1,&\text{if }s\in\{x,y\},\\ 1,&\text{otherwise}.\end{cases}$
It is easy to verify that $G$ contains a clique of size $\ell$ if and only if there is a bipartition of $V(G^{\prime})$ such that $s$ and $t$ are in different components with the one containing $s$ having $\ell+1$ vertices and, moreover, the total weight of edges crossing the bipartition is at most $n\cdot(d+1)-\frac{\ell\cdot(\ell+1)}{2}$. ∎
Based on Theorem 4 and the reduction in the proof of Theorem 1, we obtain the following result.
@@
SB-ESBP-DAG is NP-hard and is W[1]-hard with respect to $\ell$. This holds even if the function $p$ has two different values and the function $q$ has three different values.

### v3

@@ 4 Energy-Saving Partition of DAG
Let $v$ be as stipulated in Claim 6. We consider first the case where $N^{+}_{G}{[v]}\subseteq X$ (i.e., $Z=X$). In this case, $X^{\prime}=X\cup\{v^{+},v^{-}\}$ and $s\in X^{\prime}$. We need to prove that $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))\leq w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$. To this end, observe that none of the arcs of $\{(u,v^{-}):u\in N^{+}_{G}{(v)}\}$ is in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$, due to the infinite weight of these arcs (see (1b)). It follows that $v^{-}\in X$. If $v^{+}\in X$, then $(X^{\prime},Y^{\prime})=(X,Y)$, and thus $w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))=w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$; we are done. If $v^{+}\in Y$, then since $N^{-}_{G^{\prime}}{(v^{+})}=\{v\}$, we have that $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))-w({{\textsf{cut}}}_{G^{\prime}}(X^{\prime},Y^{\prime}))=w((v,v^{+}))\geq 0$.
We consider now the case where $N^{+}_{G}{[v]}\subseteq Y$ (i.e., $Z=Y$). In this case, $X^{\prime}=Y\cup\{v^{+},v^{-}\}$, $Y^{\prime}=X\setminus\{v^{+},v^{-}\}$, and $s\in Y^{\prime}$. Observe that none of the arcs in $\{(v^{+},u):u\in N^{+}_{G}{(v)}\}$ belongs to ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$, due to their infinite weight (see (1b)). It follows that $v^{+}\in Y$. If $v^{-}\in Y$, then $(Y^{\prime},X^{\prime})=(X,Y)$, and thus $w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))=w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$; we are done. If $v^{-}\in X$, then since $N^{+}_{G^{\prime}}{(v^{-})}=\{v\}$, we have that $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))-w({{\textsf{cut}}}_{G^{\prime}}(Y^{\prime},X^{\prime}))=w((v^{-},v))\geq 0$. ∎
By Claim 6, we may assume that, for every $v\in V(G)$ such that $N^{+}_{G}{(v)}\neq\emptyset$, if $N^{+}_{G}{[v]}$ are contained in the same $Z\in\{X,Y\}$, then $\{v^{+},v^{-}\}$ are also contained in $Z$.
The subsequent property essentially posits that for any vertex $v\in V(G)$ placed in part $X$, if at least one of $v$’s outneighbors in $G$ is placed in the opposite part $Y$, then the arc $(v,v^{+})$, whose weight equals the amount of energy consumption for transferring the output of $v$, must be in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$.
@@
Let $v\in X\cap V(G)$ such that $N^{+}_{G}{(v)}\cap Y\neq\emptyset$. Then, $(v,v^{+})\in{{\textsf{cut}}}_{G^{\prime}}(X,Y)$ and none of the arcs entering or leaving $v^{-}$ is contained in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$.
@@
Let $v$ be as stipulated in the claim. Let $u$ be an arbitrary vertex from $N^{+}_{G}{(v)}\cap Y$. It must hold that $v^{+}\in Y$, because if not, the arc $(v^{+},u)$, which has an infinite weight, would be included in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$, contradicting with $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$. Since $v$ is the only outneighbor of $v^{-}$ in $G^{\prime}$ and $v\in X$, the only arc $(v^{-},v)$ leaving $v^{-}$ is excluded from ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$. Finally, recall that $N^{-}_{G^{\prime}}{(v^{-})}=N^{+}_{G}{(v)}$, and by the definition of the function $w$, every arc $(u,v^{-})$ where $u\in N^{+}_{G}{(v)}$ has an infinite weight. Since $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$, none of the arcs entering $v^{-}$ can be in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$. ∎
By Claim 7, for every $v\in X\cap V(G)$ having at least one outneighbor in $G$ that is put in $Y$, we may assume that $v^{+}\in Y$. Moreover, as $v$ is the only outneighbor of $v^{-}$ in $G^{\prime}$, we may assume that $v^{-}\in X$.
Applying analogous reasoning, we derive the following assertion.
@@
Let $v\in Y\cap V(G)$ such that $N^{+}_{G}{(v)}\cap X\neq\emptyset$. Then, $(v^{-},v)\in{{\textsf{cut}}}_{G^{\prime}}(X,Y)$, and none of the arcs entering or leaving $v^{+}$ is contained in ${{\textsf{cut}}}_{G^{\prime}}(X,Y)$.
For every $v\in Y\cap V(G)$ as in Claim 8, we may assume that $v^{-}\in X$ and $v^{+}\in Y$.
Let $V_{1}=V(G)\cap X$ and let $V_{2}=V(G)\cap Y$. We show below that the assignment corresponding to $(V_{1},V_{2})$ is a Yes-witness of the instance of ESP-DAG. By the definition of the weight function $w$, we have that
@@
Let $V_{1}^{\prime}=\{v\in V_{1}:N^{+}_{G}{(v)}\cap Y\neq\emptyset\}$ and let $V_{2}=\{v\in V_{2}:N^{+}_{G}{(v)}\cap X\neq\emptyset\}$. By Claims 68, we have:
  • For every $i\in[2]$ and every $v\in V_{i}\setminus V_{i}^{\prime}$, $\{v^{+},v^{-}\}$ are contained in the same part of $(X,Y)$ as $v$, whenever $v$ is not a sink in $G$.
@@
The sum of the left sides of Equalities (9)–(11) is $\left(\sum_{v\in V_{i},i\in[2]}p(v,i)\right)+{\textsf{BTran}_{G}((V_{1},V_{2}),q)}$, and the sum of the right sides of them is $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))$. From $w({{\textsf{cut}}}_{G^{\prime}}(X,Y))\leq r$, we know that the instance of the ESP-DAG problem is a Yes-instance. ∎
As the DM-$s$-$t$-Cut problem can be solved in time $O(n\cdot m\cdot\log(n^{2}/m))$ [18], by Theorem 2, we reach the following result.
@@
ESP-DAG with $k=2$ can be solved in time $O(n\cdot m\cdot\log(n^{2}/m))$, where $n$ and $m$ are respectively the number of vertices and the number of edges of the input digraph.
Next, we derive a polynomial-time algorithm for a special case of ESP-DAG where the input DAG is a directed path. Tasks having such precedence dependencies are relevant to many applications (see, e.g., [1, 20]).
@@
The ESP-DAG problem can be solved in time $O(n\cdot k^{3}\cdot M)$ when the given DAG is a directed path with $n$ vertices, where $M$ represents the number of bits to encode the largest value of the functions $p$ and $q$.
@@
Let $I=(G,p,q,r)$ be an instance of the ESP-DAG problem, where $G$ is a directed path, $p:V(G)\times[k]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$ are two functions, and $r$ is a number. Let $M$ be the number of bits to encode the largest value of the functions $p$ and $q$. We derive a dynamic programming algorithm to solve the problem as follows.
Let $(v_{1},v_{2},\dots,v_{n})$ be the directed path representing $G$. We maintain a table $S(i,j)$, where $i\in[n]$ and $j\in[k]$. Specifically, $S(i,j)$ is defined as the value of an optimal solution to the instance $I$ restricted to the first $i$ vertices, under the condition that $v_{i}$ is assigned to machine $j$:
@@
By definition, for each $j\in[k]$, we have $S(1,j)=p(v_{1},j)$. We use the following recursion to update the table:
@@
After computing the table, we conclude that $I$ is a Yes-instance if and only if
@@
As there are at most $n\cdot k$ entries in the table, and computing each entry requires checking up to $k$ other entries and performing addition and comparison operations $O(k)$ times on numbers that can be encoded within $M$ bits, the entire table can be computed in time $O(n\cdot k^{3}\cdot M)$. ∎
@@ 5 Sized-Bounded Energy-Saving Partition of DAG with Two Machines
Now we switch our focus to the natural variant SB-ESBP-DAG, where there are only two machines and one of them is capable of executing only a limited number $\ell$ of tasks.
Recall that the SBM-$s$-$t$-Cut problem is exactly Multiway Cut where $k=2$ with an additional restriction that one of the parts in the desired bipartition contains at most $\ell$ vertices. However, despite the NP-hardness the SBM-$s$-$t$-Cut problem, the reduction in the proof of Theorem 1 is insufficient to show the NP-hardness of the SB-ESBP-DAG problem. The reason is that, according to the reduction, for a bipartiton $(V_{s},V_{t})$ of the graph $G$ in an instance of the SBM-$s$-$t$-Cut problem, in the corresponding bipartition $(V_{1},V_{2})$ of the created DAG $G^{\prime}$, $V_{1}$ needs to contain $V_{s}$ and all newly introduced vertices for vertices in $V_{s}$. However, the number of these newly introduced vertices cannot be bounded in an “exact” manner, which hinders the setup of the value of $\ell$. We thus need to modify the reduction in order to pinpoint a countable bound of $|V_{1}|$ and thus the value of $\ell$. To this end, through a reduction from the Clique problem restricted to regular graphs, we show that the SBM-$s$-$t$-Cut problem is NP-hard even in a special case where the degrees of the vertices in the input graph is nearly regular. As a matter of fact, we show that the problem is W[1]-hard with respect to $\ell$. Despite the widespread attention this problem has received, it is remarkable that such a W[1]-hardness result has not been previously documented. Therefore, our reduction bolsters the NP-hardness of SBM-$s$-$t$-Cut as examined in [9]. We believe this W[1]-hardness result is of independent interest, and hence explitecty state it in the following theroem.
A parameterized problem is a subset of $\Sigma^{*}\times\mathbb{N}$, where $\Sigma$ is a finite alphabet. A parameterized problem can be either fixed-parameter tractable (FPT) or fixed-parameter intractable. In particular, a parameterized problem is FPT if there is an algorithm which correctly determines for each instance $(I,\kappa)$ of the problem whether $(I,\kappa)$ is a Yes-instance in time $O(f(\kappa)\cdot|I|^{O(1)})$, where $f$ is a computable function and $|I|$ is the size of $I$. W[1]-hard problems are considered intractable, in the sense that unless FPT$=$W[1] (which is widely believed to be unlikely), they do not admit any FPT-algorithms. For greater details on parameterized complexity theory, we refer to [11, 14, 15].
@@
The SBM-$s$-$t$-Cut problem is W[1]-hard with respect to $\ell$. Furthermore, this holds even when the edges of the given digraph have at most two distinct weights, and all vertices, except $s$ and $t$, have the same degree.
@@
Let $(G,\ell)$ be an instance of the Clique problem, where $G$ is a regular graph with $n$ vertices. Let $d$ denote the degree of the vertices in $G$. We assume that $d>\ell$, as otherwise the problem becomes trivial to solve. Let $G^{\prime}$ be the graph obtained from $G$ by adding two new vertices, $s$ and $t$, both of which are adjacent to all vertices of $G$. Define a weight function $w:E(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ such that, for every edge $\{x,y\}\in E(G^{\prime})$, the following holds:
$w(\{x,y\})=\begin{cases}d+2,&\text{if }s\in\{x,y\},\\ 1,&\text{otherwise}.\end{cases}$
Let $r=n\cdot(d+2)-\ell^{2}$. Let $I=(G^{\prime},w,\{s,t\},r,\ell+1)$, which represents an instance of the SBM-$s$-$t$-Cut problem. Notice that both $s$ and $t$ have the same degree $n$, and every other vertex has a degree of $d+1$ in $G^{\prime}$. We prove the correctness of the reduction as follows.
()($\Rightarrow$)( ⇒ ) Assume that $G$ contains a clique $K\subseteq V(G)$ of $\ell$ vertices. Let $V_{s}=K\cup\{s\}$ and let $V_{t}=\{t\}\cup V(G)\setminus K$. Obviously, $(V_{s},V_{t})$ forms a bipartition of $V(G^{\prime})$. By the definition of the function $w$, the total weight of edges crossing $(V_{s},V_{t})$ is $(n-\ell)\cdot(d+2)+\ell\cdot(d-\ell+1)+\ell=n\cdot(d+2)-\ell^{2}=r$. Therefore, $(V_{s},V_{t})$ is a Yes-witness for the instance $I$ of the SBM-$s$-$t$-Cut problem.
()($\Leftarrow$)( ⇐ ) Assume that $I$ is a Yes-instance of the SBM-$s$-$t$-Cut problem, i.e., there exists a bipartition $(V_{s},V_{t})$ of $V(G^{\prime})$ such that $s\in V_{s}$, $|V_{s}|\leq\ell+1$, $t\in V_{t}$, and the total weight of the edges between $V_{s}$ and $V_{t}$ is at most $r$ with respect to the weight function $w$. Let $v$ be an arbitrary vertex in $V_{t}\setminus\{t\}$, if any. Observe that as the weight of the edge between $s$ and $v$ is $d+2$, and each of the other $d+1$ edges incident to $v$ in $G^{\prime}$ has a weight of $1$, moving $v$ from $V_{t}$ to $V_{s}$ strictly reduces the total weight of the edges between $V_{s}$ and $V_{t}$. Based on this observation, we may assume that $|V_{s}|=\ell+1$. Let $K=V_{s}\setminus\{s\}$, so $|K|=\ell$. The total weight of the edges between $\{s\}$ and $V(G)\setminus K$ is $(n-\ell)\cdot(d+2)$. The total weight of the edges between $\{t\}$ and $K$ is $\ell$. This implies that the total weight of the edges between $K$ and $V(G)\setminus K$ is at most $r-(n-\ell)\cdot(d+2)-\ell=\ell\cdot(d-\ell+1)$. Since each of these edges has a weight of $1$, there are at most $\ell\cdot(d-\ell+1)$ edges between $K$ and $V(G)\setminus K$ in $G$. Given that $G$ is $d$-regular, it follows that there are at least $\frac{|K|\cdot d-\ell\cdot(d-\ell+1)}{2}=\frac{\ell\cdot(\ell-1)}{2}$ edges within $K$. This is possible only if $K$ is a clique in $G$. Therefore, we can conclude that $(G,\ell)$ is a Yes-instance of the Clique problem. ∎
We are ready to show the NP-hardness of the SB-ESBP-DAG problem.
@@
The SB-ESBP-DAG problem is NP-hard. Furthermore, this holds even when the function $p$ has two distinct values and the function $q$ has three distinct values.